OpenStudy (anonymous):

Use the function to evaluate the indicated expressions and simplify. f(x) = -3x2 - 7 f(x+2)=? f(x)+f(2)=?

OpenStudy (campbell_st):

substitute (x + 2) into the function -3(x+2)^2 - 7 expand and simplify find f(2) by substituting x = 2 then simplify f(x) + f(2) = -3x^2 - 7 + (-19)

Directrix (directrix):

f( x+2)= -3(x+2)^2 - 7 =-3x^2 -12x -19 f(x) + f(2) = -3x^2 -26 f(x+2) is NOT equal to f(x) + f(2)